OpenStudy (anonymous):

Suppose a particular jet needs to attain a speed of 220 mph to take off. If it can accelerate from 0 to 220 mph in 30 seconds, how long must the runway be? Assume constant acceleration. Note: 1 mph = 22/15 ft/sec.

OpenStudy (anonymous):

so your acceleration, or a, is: a=(220mph-0)/30s but you have to change hours into seconds so 220mph=220/3600=.06111 so a=.06111/30=.002037mi/s^2 the time it takes to reach this spead is given by Vf = Vi +at or final velocity equals initial velocity plus acceleration times time so (.06111mi/s-0)/.002037mi/s^2=t=30 seconds. so the distance is x=Vit + .5at^2 or x=0+(.5)(.002037)(30)^2 = .91655 miles which equals 4839.9 feet

OpenStudy (anonymous):

4839.9 feet is the length of the runaway

OpenStudy (anonymous):

What D:< lol

OpenStudy (nincompoop):

and someone's homework was done without that person learning at all

OpenStudy (anonymous):

ok heres a better solution for him you are given: initial speed, u = 0 final speed, v = 220 acceleration = (v-u)/t ===> t= time taken for the plane to reach final velocity(in hours) then you use the formula: displacement, s = u.t + (1/2).a.t^2